# ObjectAdd: Adding Objects into Image via a Training-Free Diffusion Modification Fashion

### v2

@@ 4.1 Experimental Setup
$\mathcal{L}_{pixels}=avg(\lvert(1-\mathcal{M})\odot{\mathcal{I}}^{0}-(1-\mathcal{M})\odot\bar{\mathcal{I}}^{0}\rvert)$ (16)
This metric is to calculate the location dependent consistency of non-edited area. To measure the location independent consistency, we strictly follow [20] to calculate the histogram similarity, denoted as “Hist” in our following content.
Lastly, in order to assess the quality of the added object, we compute the text-to-image consistency using the CLIP Score as a metric:
@@
where the $clip()$ embeds the image and text into the same space, then calculats their distance. The edited images undergo masking to isolate the object. Our script for computing this metric is derived from the open-source codes [16].
4.2 Qualitative Comparison
As illustrated in Fig. 4, our ObjectAdd method outperforms all other techniques in terms of performance. Although DALL-E 3 is capable of generating high-quality images containing the desired object, it does not effectively maintain consistency between the original and edited images. Both P2P and SD-v1-4 occasionally encounter difficulties in creating the object or yield a low-quality rendition, depending on the ambiguity of the merged prompt. Moreover, they predominantly fail to uphold consistency. As a result, these qualitative results well demonstrate that our ObjectAdd is superior to the other methods.
@@ 4.2 Qualitative Comparison
4.3 Quantitative Comparison
We compute the three metrics delineated in Sec. 4.1 for all four methods to compare their quantitative outcomes. The assessment is carried out on the six pairs of prompt $\mathcal{P}$ generated images and their corresponding versions with the added word $\mathcal{W}$ described objects, as illustrated in Fig. 4. Table 1 showcases the test results, wherein our ObjectAdd method surpasses others across all metrics.
First and foremost, our ObjectAdd method demonstrates a remarkable ability to maintain the consistency of non-edited areas by achieving a reduction of at least 20 “By Pixels” compared to other techniques. Secondly, in terms of the “Hist” metric, DALL-E 3 and SD-v1-4 trail significantly, while P2P displays a performance closely resembling ours. This can be attributed to the minimal alterations made to the image by P2P, as depicted in Fig. 4. Nonetheless, our proposed ObjectAdd method remains the frontrunner. Lastly, when examining the CLIP score, the three comparison methods exhibit similar performance, whereas our ObjectAdd method significantly outperforms them all. This suggests that our approach is more adept at generating the desired object in the appropriate location. These observations are congruent with the visual findings in Fig. 4.
4.4 Ablation Study
We engage in ablation studies to scrutinize the effectiveness of each module introduced in this paper. These include text embedding coalesce, object-driven layout control (incorporating both latent and attention information injection), and prompted image inpainting. We confirm their efficacy by sequentially integrating each module and visually presenting the outputs in Fig. 6.111More ablations w.r.t. time steps to incorporate object-driven layout control and prompted image inpainting have been provided in the Appendix.
Table 1: A comparison of quantitative results.
By Pixels ($\downarrow$) Hist ($\uparrow$) CLIP Score ($\uparrow$)
DALL-E 3 117.126 0.710 22.878
P2P 115.192 0.981 22.709
SD-v1-4 127.572 0.792 22.394
ObjectAdd (Ours) 99.139 0.987 23.180
Fig. 6 clearly demonstrates that the omission of these methodologies leads to outputs that bear no relevance to the prompts. This issue arises due to our approach’s reliance on text embedding coalescence for the generation of image embeddings. In the absence of this process, there is no conditioned input for generation, resulting in arbitrary images. As depicted in Fig. 6, text embedding coalescence facilitates the emergence of principal subjects, such as a woman and a hat. However, without layout control, the output suffers from a lack of consistency. This challenge is largely mitigated by the introduction of latent injection, causing the results to more closely resemble our final output. Yet, the details of the hat appear distorted, exhibiting two brims as shown in Fig. 6. This issue is rectified through the application of attention injection, which enhances the correlation between the mask area and the object prompt, as evidenced in Fig. 6. The final obstacle pertains to the background. As illustrated in Fig. 6, we employ prompted image inpainting to achieve non-edited area consistency.
(j) None
(k) + TEC
(l) + LII
(m) + AII
(n) + PII
@@ 4.4 Ablation Study
4.5 Limitation and Future Work
Moving forward, the focus will be on developing a refined approach that ensures seamless accessibility for non-specialists. Besides, exploring advanced open models like SD-v3 version [25] is crucial. Finally, creating a suite of automated evaluation tools for comprehensive assessment is of paramount importance.
5 Conclusion
In this study, we have presented a groundbreaking training-free methodology for incorporating objects into a user-defined area of a diffusion image while preserving the remainder of the image with minimal alterations. As a pioneering effort in this domain, our proposed ObjectAdd technique employs embedding-level concatenation to guarantee accurate text embedding coalescence, object-driven layout control through latent and attention injection to ensure objects are placed within the user-specified area, and prompted image inpainting utilizing attention refocusing and object expansion strategies to maintain the integrity of the rest of the image. A thorough combination of qualitative and quantitative comparisons with closely related works, such as DALL-E 3, P2P, and SD-v1-4, have convincingly showcased the efficacy of our innovative ObjectAdd methodology.
Appendix 0.A More Ablations
Herein, we delve deeper into our research by offering additional ablative studies concerning time steps for object-driven layout control (latent information injection and attention information injection), and the specific step for prompted image inpainting. Our comprehensive analysis aims to shed more light on the intricate dynamics of these processes and provide a more nuanced understanding of their roles in image generation and manipulation.
@@ Appendix 0.A More Ablations
0.A.0.1 Latent Information Injection.
In Fig. 7, we present the outcomes of various latent injection steps, demonstrating the effectiveness of this approach in generating high-quality images. As the time steps $t$ approach zero from $T$, we observe that the optimal performance is achieved when the latent information is injected during the initial 20% of the time steps. Employing an excessive number of steps for this operation may cause the appearance of unnatural edges and inconsistencies within the object, whereas utilizing too few steps might compromise the overall quality of the generated image.
@@
0.A.0.2 Attention Information Injection.
As shown in Fig. 8, similar to what occurs in latent information injection, too many steps of injection cause the object to disappear, while fewer steps than our setting can lead to poor quality. The most appropriate steps falls into the first 30% diffusion steps.
@@
0.A.0.3 Prompted Image Inpainting.
In the main paper, we have discussed the optimal selection of the time step $t$ for prompted image inpainting, suggesting that it should fall within the range of $0.3\cdot T<t<0.5\cdot T$. Our experimental findings, as depicted in Fig. 9, substantiate this claim. Images (b) through (d) exhibit remarkable similarity, indicating that variations in the chosen $t$ values do not significantly impact the final outcomes. Conversely, images (a) and (e) demonstrate inferior performance, signifying that selecting $t$ values beyond the recommended range is ill-advised.

### v5

@@ IV-A3 Evaluation Metrics
$\mathcal{L}_{pixels}=avg(\lvert(1-\mathcal{M})\odot{\mathcal{I}}^{0}-(1-\mathcal{M})\odot\bar{\mathcal{I}}^{0}\rvert),$ (18)
which calculates the consistency of non-edited area.
Then, in order to assess the quality of the added object, we compute the text-to-image consistency using the CLIP Score as a metric:
@@
where the $clip()$ embeds the image and text into the same space, then calculates their distance. The edited images undergo masking to isolate the object. Our script for computing this metric is derived from the open-source codes [27].
We also apply FID [28] metric, since it is a commonly used one for measuring image generation quality.
IV-A4 Data Preparation
To facilitate quantitative comparison, we have created our own dataset. The first step involves generating prompts. We provide a few example prompts such as “A woman wearing glasses” and “A lake” to ChatGPT and request it to generate 320 similar prompts. Using these prompts, we then generate corresponding images. During this step, we fix the random seeds to ensure reproducibility, allowing us to regenerate the same images given the same prompts.
Following image generation, we use a custom script to draw bounding boxes on these images. The box information for each image is saved in a text file with five lines, detailing the coordinates of the top-left corner, the width and height of the box, and the prompt used to add the object in the final line. In cases where the original image is of low quality, we discard it, resulting in a final dataset of 229 images.
Our model reads these text files according to the prompts, generates the same original image $\mathcal{I}$, and uses the information in the text file to create the mask $\mathcal{M}$, subsequently generating the reference image $\tilde{\mathcal{I}}$.
(a) A man wearing glasses
A hat
(b) A young woman wearing glasses
A beret
(c) A boy wearing glasses
A cap
(d) A lake
A plane
(e) A lake
A warship
(f) A lake
A whale
(g) A whole round table
A basketball
(h) A whole round table
A pumpkin
(i) A whole round table
A orange
IV-B Qualitative Comparison
As illustrated in Fig. 5, our ObjectAdd method outperforms all other techniques in terms of performance. InstructPix2Pix fails to generate the adding object in most cases, while been unnatural and inconsistent in the case of apples. Both P2P and SD-v1-4 occasionally encounter difficulties in creating the object or yield a low-quality rendition, depending on the ambiguity of the merged prompt. Moreover, they predominantly fail to uphold consistency. As a result, these qualitative results well demonstrate that our ObjectAdd is superior to the other methods.
@@ IV-B Qualitative Comparison
(a) None
(b) + TEC
(c) + LII
(d) + AII
(e) + PII
@@
TABLE I: A comparison of quantitative results.
By Pixels ($\downarrow$) FID ($\downarrow$) CLIP Score ($\uparrow$)
InstructPix2Pix 11.393 39.559 18.129
P2P 12.833 46.106 17.809
SD-v1-4 49.824 157.821 18.724
ObjectAdd (Ours) 10.219 82.437 19.424
IV-C Quantitative Comparison
We compute the three metrics delineated in Sec. IV-A for all four methods to compare their quantitative outcomes. The assessment is carried out on the 229 pairs of prompts $\mathcal{P}$ generated images and their corresponding versions with the added word $\mathcal{W}$ described objects, as well as the box that illustrates the required position for object generation. Table I showcases the test results, wherein our ObjectAdd method surpasses others in general.
First and foremost, our ObjectAdd method demonstrates a remarkable ability to maintain the consistency of non-edited areas, achieving the best performance in “By Pixels” compared to other techniques. Additionally, in terms of the “CLIP Score” metric, our approach significantly surpasses all other methods. As mentioned, “By Pixels” indicates the consistency between the non-masked areas of the edited image and the original one, while the “CLIP Score” measures the extent to which the masked area fits the corresponding prompt. Thus, the results of these two metrics suggest that our approach performs the best on both masked and unmasked areas compared to other methods.
The “FID” score of our method does not seem satisfactory; however, this is because other methods barely edit the original image. Since FID measures the difference in the distribution between two datasets—original images and edited images—a lower value suggests better consistency between the original and edited images. However, the “By Pixels” metric indicates that our approach has the best consistency in the unmasked areas. Therefore, we can conclude that the better FID scores of InstructPix2Pix and P2P compared to our method are due to their limited image editing capabilities, which is consistent with the “CLIP Score” metric.
IV-D Ablation Study
We engage in ablation studies to scrutinize the effectiveness of each module introduced in this paper. These include text embedding coalesce, object-driven layout control (incorporating both latent and attention information injection), and prompted image inpainting. We confirm their efficacy by sequentially integrating each module and visually presenting the outputs in Fig. 7.
Fig. 7(a) clearly demonstrates that the omission of these methodologies leads to outputs that bear no relevance to the prompts. This issue arises due to our approach’s reliance on text embedding coalescence for the generation of image embeddings. In the absence of this process, there is no conditioned input for generation, resulting in arbitrary images. As depicted in Fig. 7(b), text embedding coalescence facilitates the emergence of principal subjects, such as a woman and a hat. However, without layout control, the output suffers from a lack of consistency. This challenge is largely mitigated by the introduction of latent injection, causing the results to more closely resemble our final output. Yet, the details of the hat appear distorted, exhibiting two brims as shown in Fig. 7(c). This issue is rectified through the application of attention injection, which enhances the correlation between the mask area and the object prompt, as evidenced in Fig. 7(d). The final obstacle pertains to the background. As illustrated in Fig. 7(e), we employ prompted image inpainting to achieve non-edited area consistency.
Further more, we delve deeper into our research by offering additional ablative studies concerning time steps for object-driven layout control (latent information injection and attention information injection), and the specific step for prompted image inpainting. Our comprehensive analysis aims to shed more light on the intricate dynamics of these processes and provide a more nuanced understanding of their roles in image generation and manipulation.
@@ IV-D Ablation Study
Latent Information Injection. In Fig. 8, we present the outcomes of various latent injection steps, showing the efficacy of this approach in generating high-quality images. As the time steps $t=T\rightarrow 0$, we observe that the optimal performance is achieved when the latent information is injected during the initial 20% of the time steps. Employing an excessive number of steps for this operation may cause the appearance of unnatural edges and inconsistencies within the object, whereas utilizing too few steps might compromise the overall quality of the generated image.
@@
Attention Information Injection. As shown in Fig. 9, similar to what occurs in latent information injection, too many steps of injection cause the object to disappear, while fewer steps than our setting can lead to poor quality. The most appropriate steps falls into the first 30% steps.
@@
Prompted Image Inpainting. In previous sections, we have discussed the optimal selection of the time step $t$ for prompted image inpainting, suggesting that it should fall within the range of $0.3\cdot T<t<0.5\cdot T$. Our experimental findings, as depicted in Fig. 10, substantiate this claim. Images (b) through (d) exhibit remarkable similarity, indicating that variations in the chosen $t$ values do not significantly impact the final outcomes. Conversely, images (a) and (e) demonstrate inferior performance, signifying that selecting $t$ values beyond the recommended range is ill-advised.
@@
Figure 11: Example of inserting user specified images.
IV-E Performance with a Real Image as Input
As mentioned in previous sections, our approach also allows users to take an image as input for inserting it into the boxing area. In this section, we will display some results of this function and discuss them. As illustrated in Fig. 11, our approach successfully generates the required object in the boxing area while preserving the main features of the input image, such as the color of boats and the texture of top hats. One interesting observation is that the prompt for the inserted image should accurately reflect its content; otherwise, the result may be of poor quality after insertion.
IV-F Limitation and Future Work
Moving forward, the focus will be on developing a refined approach that ensures seamless accessibility for non-specialists. Besides, exploring advanced open models like SD-v3 version [29] is crucial. Finally, creating a suite of automated evaluation tools for comprehensive assessment is of paramount importance. Besides, it would be an interesting idea to transform input image to prompt automatically using a pre-trained model like BLIP [30], which might make our method more practical for using.
V Conclusion
In this study, we have presented a groundbreaking training-free methodology for incorporating objects into a user-defined area of a diffusion image while preserving the remainder of the image with minimal alterations. As a pioneering effort in this domain, our proposed ObjectAdd technique employs embedding-level concatenation to guarantee accurate text embedding coalescence, object-driven layout control through latent and attention injection to ensure objects are placed within the user-specified area, and prompted image inpainting utilizing attention refocusing and object expansion strategies to maintain the integrity of the rest of the image. A thorough combination of qualitative and quantitative comparisons with closely related works, such as DALL-E 3, P2P, InstructPix2Pix, and SD-v1-4, have convincingly showcased the efficacy of our innovative ObjectAdd methodology.
Acknowledgments
This work was supported by National Science and Technology Major Project (No. 2022ZD0118202), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).